# Ring Attention with Blockwise Transformers for Near-Infinite Context

### v1

@@ 5.2 Evaluating Model Flops Utilization
5.3 Impact on LLM Performance
We evaluate Ring Attention by applying our method to finetune LLaMA model to longer context. In this experiment, while our approach enables training with millions of context tokens, we conducted finetuning on the LLaMA-13B model, limiting the context length to 512K tokens due to constraints on our cloud compute budget. This finetuning was carried out on 8 A100 GPUs, using the ShareGPT dataset, following methodologies as outlined in prior works (Chiang et al., 2023; Geng et al., 2023). We then evaluated our finetuned model on the line retrieval test (Li et al., 2023). In this test, the model needs to precisely retrieve a number from a long document, the task can effectively capture the abilities of text generation, retrieval, and information association at long context, reflected by the retrieving accuracy. Figure 3 presents the accuracy results for different models across varying context lengths (measured in tokens). Notably, our model, Ring Attention-13B-512K, stands out as it maintains high accuracy levels even with long contexts. GPT3.5-turbo-16K, Vicuna-16B-16K, and Claude-2-100K demonstrate competitive accuracy within short context lengths. However, they cannot handle extended context lengths.
6 Related Work
Transformers have garnered significant attention in the field of AI and have become the backbone for numerous state-of-the-art models. Several works have explored memory-efficient techniques to address the memory limitations of Transformers and enable their application to a wider range of problems. Computing exact self-attention in a blockwise manner using the tiling technique (Milakov and Gimelshein, 2018) has led to the development of memory efficient attention mechanisms (Rabe and Staats, 2021) and its efficient CUDA implementation (Dao et al., 2022), and blockwise parallel transformer (Liu and Abbeel, 2023b) that proposes computing both feedforward and self-attention block-by-block, resulting in a significant reduction in memory requirements. In line with these advancements, our work falls into the category of memory efficient computation for Transformers. Other works have investigated the approximation of attention mechanisms, yet these efforts have often yielded sub-optimal results or encountered challenges during scaling up. For an in-depth review of these techniques, we recommend referring to the surveys (Narang et al., 2021; Tay et al., 2022). Another avenue of research explores various parallelism methods, including data parallelism (Dean et al., 2012), tensor parallelism (Shoeybi et al., 2019), pipeline parallelism (Narayanan et al., 2019; Huang et al., 2019; Narayanan et al., 2021), sequence parallelism (Li et al., 2021; Korthikanti et al., 2022; Jacobs et al., 2023), and FSDP (Facebook, 2023; Rajbhandari et al., 2020). The activations of self-attention take a substantial amount of memory for large context models. Tensor parallelism can only reduce parts of activations memory and sequence parallelism introduces a significant communication overhead that cannot be fully overlapped with computation. Our work leverages on blockwise parallel transformers to distribute blockwise attention and feedforward across devices and concurrently overlaps the communication of key-value blocks in a circular of hosts with the computation of query-key-value blocks and feedforward, allowing number of devices times longer sequences with negligible overheads. Overlapping communication with computation has been studied in high performance computing literature (Danalis et al., 2005; Wang et al., 2022; Danalis et al., 2009, inter alia). While ring communication has found applications in other parallel computing scenarios (Bischof, 2008; Hursey and Graham, 2011; Gibiansky, 2017; Sergeev and Del Balso, 2018), our work stands out as the first work to show that it can be applied to self-attention as used in Transformers and to make it fit efficiently into Transformer training and inference without adding significant overhead by overlapping blockwise computation and communication.
7 Conclusion
In conclusion, we propose a memory efficient approach to reduce the memory requirements of Transformers, the backbone of state-of-the-art AI models. Our approach allows the context length to scale linearly with the number of devices while maintaining performance, eliminating the memory bottleneck imposed by individual devices. Through extensive experiments on language modeling and reinforcement learning, we demonstrate its effectiveness, enabling training sequences that are device count times longer than those of prior memory-efficient Transformers, exceeding a context length of 100 million without making approximations to attention.
Limitations and Future Work. Although our method achieves state-of-the-art context length for Transformer models, it does have some limitations that need to be addressed:
  • Scaled up training: due to compute budget constraint, our experiments focus on evaluation the effectiveness of the proposed approach without large scale training models.
  • Optimal compute performance: While Ring Attention scales context length linearly with device count while maintaining performance, optimizing low-level operations is required for achieving optimal compute performance. We suggest considering porting our method to CUDA, OpenAI Triton, or Jax Pallas in the future, for both maximum sequence length and maximum compute performance.
In terms of future prospects, the possibility of near-infinite context introduces a vast array of exciting opportunities, such as large video-language models, decision making and tool using transformers on extended trial-and-error experience, training models on large codebases, and adapting AI models to understand science such as gene sequences.
Acknowledgments
Appendix A Code
The implementation of Ring Attention in Jax is provided in Figure 4. We use defvjp function to define both the forward and backward passes, and use collective operation jax.lax.ppermute to facilitate the exchange of key-value blocks among a ring of hosts. The provided code snippet highlights essential components of Ring Attention. The complete implementation with maximum memory efficient just needs to replace the local blockwise computation, specifically jnp.einsum("bshd,btd->bhst", q, k) and jnp.einsum("bhst,btd->bshd", s, v) as well as the local blockwise feedforward computation with BPT’s Jax based blockwise attention and FFN computation [^2] https://github.com/lhao499/llm_large_context. For maximum compute efficiency our Ring Attention can be integrated with exiting kernel-level fused-attention implementations, such as on GPUs Ring Attention can be integrated with Triton code [kernel team, 2023] or CUDA code [Dao et al., 2022], similarly on TPUs it is also compatible with Pallas code [team, 2023] of the memory efficient attention [Rabe and Staats, 2021].
@partial(jax.custom_vjp, nondiff_argnums=[3, 4, 5])
def _ring_attention_fwd(q, k, v, mask, axis_name, float32_logits):
if float32_logits:
q, k = q.astype(jnp.float32), k.astype(jnp.float32)
batch, q_len, num_heads, _ = q.shape
%**** main.tex Line 450 ****batch, kv_len, dim_per_head = k.shape
axis_size = lax.psum(1, axis_name)
scale = jnp.sqrt(q.shape[-1])
def scan_kv_block(carry, idx):
prev_max_score, numerator, denominator, k, v = carry
mask = lax.dynamic_slice_in_dim(mask,
(lax.axis_index(axis_name) - idx) attn_weights = jnp.einsum("bqhd,bkd->bhqk", q, k) / scale
attn_weights = jnp.where(mask, -jnp.inf, attn_weights)
max_score = jnp.maximum(prev_max_score, jnp.max(attn_weights, axis=-1))
exp_weights = jnp.exp(attn_weights - max_score[…, None])
correction = rearrange(jnp.exp(prev_max_score - max_score), b h q -> b q h’)[…, None]
numerator = numerator * correction + jnp.einsum("bhqk,bkd->bqhd", exp_weights, v)
denominator = denominator * jnp.exp(prev_max_score - max_score) + jnp.sum(exp_weights, axis=-1)
k, v = map(lambda x: lax.ppermute(x, axis_name, perm=[(i,
(i + 1) return (max_score, numerator, denominator, k, v), None
(numerator, max_score, denominator, _, _), _ = lax.scan(scan_kv_block,
output = numerator / rearrange(denominator, b h q -> b q h’)[…, None]
return output.astype(v.dtype), (output, q, k, v, numerator, denominator, max_score)
[…]
del float32_logits
axis_size = lax.psum(1, axis_name)
output, q, k, v, numerator, denominator, max_score = res
dq = jnp.zeros_like(q, dtype=jnp.float32)
dk = jnp.zeros_like(k, dtype=jnp.float32)
dv = jnp.zeros_like(v, dtype=jnp.float32)
batch, kv_len, dim_per_head = k.shape
scale = jnp.sqrt(q.shape[-1])
def scan_kv_block(carry, idx):
dq, dk, dv, k, v = carry
mask = lax.dynamic_slice_in_dim(mask,
(lax.axis_index(axis_name) - idx) attn_weights = jnp.einsum("bqhd,bkd->bhqk", q, k) / scale
attn_weights = jnp.where(mask, -jnp.inf, attn_weights)
exp_weights = jnp.exp(attn_weights - max_score[…, None]) / denominator[…, None]
ds = jnp.einsum("bqhd,bkd->bhqk", g, v)
dl = (ds - jnp.einsum("bqhd,bqhd->bhs", g, output)[…, None]) * exp_weights
dq = dq + jnp.einsum("bhqk,bkd->bqhd", dl, k) / scale
dk = dk + jnp.einsum("bqhd,bhqk->bkd", q, dl) / scale
dv = dv + jnp.einsum("bhqk,bqhd->bkd", exp_weights, g)
k, v, dk, dv = map(lambda x: lax.ppermute(x, axis_name, perm=[(i,
(i + 1) return (dq, dk, dv, k, v), None
%**** main.tex Line 500 ****dq, dk, dv = dq.astype(q.dtype), dk.astype(k.dtype), dv.astype(v.dtype)
return dq, dk, dv
[…]
def ring_attention(q, k, v, mask, axis_name, float32_logits=True):
y, _ = _ring_attention_fwd(q, k, v, mask, axis_name, float32_logits)
return y
[…]
Appendix B Experiment Details
B.1 Evaluation of context length
In the experimental results presented in Section 5.1, we used tensor parallelism to partition the model across GPUs or TPU units. Our evaluation focused on determining the maximum achievable sequence length, using a sequence number of one. For TPUs, we utilized its default training configuration, which involved performing matmul operations in bfloat16 format with weight accumulation in float32. On the other hand, for GPUs, we adopted the default setup, where all operations were performed in float32.
B.2 Evaluation of MFU
In the evaluation presented in Section 5.2, the training was conducted using FSDP Facebook [2023] with no gradient accumulation. The batch size in tokens is 2 million per batch on GPU and 4 million per batch on TPU. For gradient checkpointing [Chen et al., 2016], we used nothing_saveable as checkpointing policies for attention and feedforward network (FFN). For more details, please refer to Jax documentation.
B.3 Evaluation on line retrieval
In the evaluation presented in Section 5.3, the training was conducted using FSDP on 8x A100 80GB Cloud GPUs. We finetuned the LLaMA-13B model [Touvron et al., 2023], limiting context length to 512K tokens due to constraints on our cloud compute budget, though our approach enables training with millions of context tokens. We use user-shared conversations gathered from ShareGPT.com with its public APIs for finetuning, following methodologies as outlined in prior works [Chiang et al., 2023, Geng et al., 2023]. ShareGPT is a website where users can share their ChatGPT conversations. To ensure data quality, we convert the HTML back to markdown and filter out some inappropriate or low-quality samples, which results in 125K conversations after data cleaning.
Appendix C Training FLOPs Scaling of Context Size
Given that our proposed approach unlocks the possibility of training with a context size exceeding 100 million tokens and allows for linear scaling of the context size based on the number of devices, it is essential to understand how the training FLOPs per dataset scale with the context size. While a larger context size results in a higher number of FLOPs, the increased ratio does not scale quadratically because the number of tokens remains fixed. We present these results in Figure 5, which showcases various model sizes and context lengths, representing different computational budgets. The figure illustrates the ratio of FLOPs for larger context lengths compared to the same model with a shorter 4K context size. We calculated the per sequence FLOPs using $(24bsh^{2}+4bs^{2}h)n$ where $h$ is model hidden dimension, $b$ is batch size, $s$ is total sequence length, and $n$ is number of layers. The per dataset FLOPs ratio is then given by $((24bs_{2}h^{2}+4b{s_{2}}^{2}h)/(24bs_{1}h^{2}+4b{s_{1}}^{2}h))/(s_{2}/s_{1})=(6h+s_{2})/(6h+s_{1})$, where $s_{2}$ and $s_{1}$ are new and old context lengths. Model sizes and their hidden dimensions are as follows: LLaMA-7B (4096), LLaMA-13B (5140), LLaMA-33B (7168), LLaMA-65B (8192), GPT3-175B (12288), and 1TB (36864). These model configurations are from LLaMA [Touvron et al., 2023] and GPT-3 [Brown et al., 2020] papers, except the 1TB model size and dimension were defined by us.
Appendix D Impact on In Context RL Performance
In addition to show the application of Ring Attention to finetune LLM in Section 5.3, we present additional results of applying Ring Attention for learning trial-and-error RL experience using Transformers. We report our results in Table 5, where we evaluate our proposed model on the ExoRL benchmark across six different tasks. On ExoRL, we report the cumulative return, as per ExoRL [Yarats et al., 2022]. We compare BC, DT [Chen et al., 2021], AT [Liu and Abbeel, 2023a], and AT with memory efficient attention [Rabe and Staats, 2021] (AT+ME), AT with blockwise parallel transformers [Liu and Abbeel, 2023b] (AT+BPT), and AT with our Ring Attention (AT+Ring Attention). The numbers of BC, DT, AT are from the ExoRL and AT paper. AT + Ring Attention numbers are run by ourselves. Since the ExoRL data is highly diverse, having been collected using unsupervised RL [Laskin et al., 2021], it has been found that TD learning performs best, while behavior cloning struggles [Yarats et al., 2022]. AT [Liu and Abbeel, 2023a] shows that conditioning Transformer on multiple trajectories with relabeled target return can achieve competitive results with TD learning. For more details, please refer to their papers. We are interested in applying Ring Attention to improve the performance of AT by conditioning on a larger number of trajectories rather than 32 trajectories in prior works. It is worth noting that each trajectory has $1000\times 4$ length where $1000$ is sequence length while $4$ is return-state-action-reward, making training 128 trajectories with modest 350M size model infeasible for prior state-of-the-art blockwise parallel transformers. Results in Table 5 show that, by scaling up the sequence length (number of trajectories), AT + Ring Attention consistently outperforms oringal AT with BPT across all six tasks, achieving a total average return of 113.66 compared to the AT with BPT model’s total average return of 111.13. The results show that the advantage of Ring Attention for training and inference with long sequences.

### v4

@@ 5.2 Evaluating Model Flops Utilization
5.3 Impact on In Context RL Performance
We present results of applying Ring Attention for learning trial-and-error RL experience using Transformers. We report our results in Table 5, where we evaluate our proposed model on the ExoRL benchmark across six different tasks. On ExoRL, we report the cumulative return, as per ExoRL (Yarats et al., 2022). We compare BC, DT (Chen et al., 2021), AT (Liu and Abbeel, 2023a), and AT with memory efficient attention (Rabe and Staats, 2021) (AT+ME), AT with blockwise parallel transformers (Liu and Abbeel, 2023b) (AT+BPT), and AT with our Ring Attention (AT+Ring Attention). The numbers of BC, DT, AT are from the ExoRL and AT paper. AT + Ring Attention numbers are run by ourselves. Since the ExoRL data is highly diverse, having been collected using unsupervised RL (Laskin et al., 2021), it has been found that TD learning performs best, while behavior cloning struggles (Yarats et al., 2022). AT (Liu and Abbeel, 2023a) shows that conditioning Transformer on multiple trajectories with relabeled target return can achieve competitive results with TD learning. For more details, please refer to their papers.
@@ 5.3 Impact on In Context RL Performance
We are interested in applying Ring Attention to improve the performance of AT by conditioning on a larger number of trajectories rather than 32 trajectories in prior works. It is worth noting that each trajectory has $1000\times 4$ length where $1000$ is sequence length while $4$ is return-state-action-reward, making training 128 trajectories with modest 350M size model infeasible for prior state-of-the-art blockwise parallel transformers. Results in Table 5 show that, by scaling up the sequence length (number of trajectories), AT + Ring Attention consistently outperforms oringal AT with BPT across all six tasks, achieving a total average return of 113.66 compared to the AT with BPT model’s total average return of 111.13. The results show that the advantage of Ring Attention for training and inference with long sequences.
5.4 Impact on LLM Performance
We evaluate Ring Attention by applying our method to finetune LLaMA model to longer context. In this experiment, while our approach enables training with millions of context tokens, we conducted finetuning on the LLaMA-13B model, limiting the context length to 512K tokens due to constraints on our cloud compute budget. This finetuning was carried out on 32 A100 GPUs, using the ShareGPT dataset, following methodologies as outlined in prior works (Chiang et al., 2023; Geng et al., 2023). We then evaluated our finetuned model on the line retrieval test (Li et al., 2023a). In this test, the model needs to precisely retrieve a number from a long document, the task can effectively capture the abilities of text generation, retrieval, and information association at long context, reflected by the retrieving accuracy. Figure 3 presents the accuracy results for different models across varying context lengths (measured in tokens). Notably, our model, Ring Attention-13B-512K, stands out as it maintains high accuracy levels even with long contexts. GPT3.5-turbo-16K, Vicuna-16B-16K, and Claude-2-100K demonstrate competitive accuracy within short context lengths. However, they cannot handle extended context lengths.
6 Related Work
Transformers have garnered significant attention in the field of AI and have become the backbone for numerous state-of-the-art models. Several works have explored memory-efficient techniques to address the memory limitations of Transformers and enable their application to a wider range of problems. Computing exact self-attention in a blockwise manner using the tiling technique (Milakov and Gimelshein, 2018) has led to the development of memory efficient attention mechanisms (Rabe and Staats, 2021) and its efficient CUDA implementation (Dao et al., 2022), and blockwise parallel transformer (Liu and Abbeel, 2023b) that proposes computing both feedforward and self-attention block-by-block, resulting in a significant reduction in memory requirements. In line with these advancements, our work falls into the category of memory efficient computation for Transformers. Other works have investigated the approximation of attention mechanisms, yet these efforts have often yielded sub-optimal results or encountered challenges during scaling up. For an in-depth review of these techniques, we recommend referring to the surveys (Narang et al., 2021; Tay et al., 2022). Another avenue of research explores various parallelism methods, including data parallelism (Dean et al., 2012), tensor parallelism (Shoeybi et al., 2019), pipeline parallelism (Narayanan et al., 2019; Huang et al., 2019; Narayanan et al., 2021), sequence parallelism (Li et al., 2023b; Korthikanti et al., 2022; Jacobs et al., 2023), and FSDP (Facebook, 2023; Rajbhandari et al., 2020). The activations of self-attention take a substantial amount of memory for large context models. Tensor parallelism can only reduce parts of activations memory and sequence parallelism introduces a significant communication overhead that cannot be fully overlapped with computation. Prior work has studied sharding along sequence and attention heads, and gathering sequences via an optimized all-to-all topology, achieving reduced communication (Jacobs et al., 2023). However, this method is restricted by the number of attention heads and requires gathering the full sequence on each device. In comparison, our approach fully overlaps communication with blockwise computation, enhancing its scalability. Prior work extends sequence parallelism for computing self-attention using a ring topology (Li et al., 2023b), which reduces the communication cost compared to standard sequence parallelism. However, overlapping communication with computation remains challenging due to the constraints of arithmetic intensity. The communication overheads render this approach infeasible for training and inference in large-context scenarios. Our work leverages on blockwise parallel transformers to distribute blockwise attention and feedforward across devices and concurrently overlaps the communication of key-value blocks in a circular of hosts with the computation of query-key-value blocks and feedforward, reducing memory cost substantially and allowing device count times larger context size with zero overheads. Overlapping communication with computation has been studied in high performance computing literature (Danalis et al., 2005; Wang et al., 2022; Danalis et al., 2009, inter alia). While ring communication has found applications in other parallel computing scenarios (Bischof, 2008; Hursey and Graham, 2011; Gibiansky, 2017; Sergeev and Del Balso, 2018), our work stands out as the first work to show that it can be applied to self-attention as used in Transformers and to make it fit efficiently into Transformer training and inference without adding significant overhead by overlapping blockwise computation and communication.
7 Conclusion
In conclusion, we propose a memory efficient approach to reduce the memory requirements of Transformers, the backbone of state-of-the-art AI models. Our approach allows the context length to scale linearly with the number of devices while maintaining performance, eliminating the memory bottleneck imposed by individual devices. Through extensive experiments on language modeling and reinforcement learning, we demonstrate its effectiveness, enabling training sequences that are up to device count times longer than those of prior memory-efficient Transformers, exceeding a context length of 100 million without making approximations to attention. In terms of future prospects, the possibility of near-infinite context introduces a vast array of exciting opportunities, such as large video-audio-language models, learning from extended feedback and trial-and-errors, understanding and generating codebase, adapting AI models to understand scientific data such as gene sequences, and developing strong reasoning from link gathering data.
Acknowledgments
Appendix A Code
The implementation of Ring Attention in Jax is provided in Figure 4. We use defvjp function to define both the forward and backward passes, and use collective operation jax.lax.ppermute to facilitate the exchange of key-value blocks among a ring of hosts. The provided code snippet highlights essential components of Ring Attention. We provide the complete code of our Ring Attention at https://github.com/lhao499/llm_large_context.
For large scale end-to-end training on TPU or on GPU cluster with high bandwidth inter connection, we recommend using FSDP to shard large models and using Ring Attention to achieve large context. If total batch size is too large, add tensor parallelism to reduce the global batch size. The degree of parallelism can be adjusted using the mesh_dim parameter within the codebase. To illustrate, consider a setup with 512 devices, such as 512x A100. If the model size is 30B, you can shard it across 8 devices and allocate the remaining 32 devices for Ring Attention. This setup allows the context size to be expanded 32 times more than if you didn’t use Ring Attention. Conversely, for models sized 7B or 3B, there is no need for FSDP. This means you can utilize all 512 devices exclusively to expand the context using Ring Attention by 512 times. Building upon the result that our approach allows for a 256K context size when using 8x A100 GPUs, it suggests that by employing 512 A100 GPUs, the potential context size can be expanded to 16 million.
def _ring_attention_fwd(q, k, v, attn_bias, axis_name, float32_logits, blockwise_kwargs):
if float32_logits:
q, k = q.astype(jnp.float32), k.astype(jnp.float32)
batch, q_len, num_heads, dim_per_head = q.shape
batch, kv_len, num_heads, dim_per_head = k.shape
axis_size = lax.psum(1, axis_name)
block_size = q_len # assumes this function is pre-sharded inside shard_map
query_chunk_size = blockwise_kwargs["query_chunk_size"]
key_chunk_size = blockwise_kwargs["key_chunk_size"]
def scan_kv_block(carry, idx):
prev_max_score, numerator, denominator, k, v = carry
attn_bias_slice = lax.dynamic_slice_in_dim(attn_bias,
(lax.axis_index(axis_name) - idx) %**** main.tex Line 500 ****q_block_idx = lax.axis_index(axis_name)
k_block_idx = (lax.axis_index(axis_name) - idx) q_chunk_idx_start = q_block_idx * (block_size // query_chunk_size)
k_chunk_idx_start = k_block_idx * (block_size // key_chunk_size)
numerator, denominator, max_score = _blockwise_attention_fwd(q, k, v,
(numerator, denominator, prev_max_score), q_chunk_idx_start, k_chunk_idx_start,
bias=attn_bias_slice, **blockwise_kwargs)
k, v = map(lambda x: lax.ppermute(x, axis_name, perm=[(i, (i + 1) for i in range(axis_size)]), (k, v))
return (max_score, numerator, denominator, k, v), None
(max_score, numerator, denominator, _, _), _ = lax.scan(scan_kv_block,
output = numerator / rearrange(denominator, b h q -> b q h’)[…, None]
return output.astype(v.dtype), (output, q, k, v, attn_bias, denominator, max_score)
[…]
output, q, k, v, attn_bias, denominator, max_score = res
batch, kv_len, num_heads, dim_per_head = k.shape
axis_size = lax.psum(1, axis_name)
dq = jnp.zeros_like(q, dtype=jnp.float32)
dk = jnp.zeros_like(k, dtype=jnp.float32)
dv = jnp.zeros_like(v, dtype=jnp.float32)
query_chunk_size = blockwise_kwargs["query_chunk_size"]
key_chunk_size = blockwise_kwargs["key_chunk_size"]
%**** main.tex Line 525 ****block_size = q.shape[1] # assumes this function is pre-sharded inside shard_map
def scan_kv_block(carry, idx):
dq, dk, dv, k, v = carry
attn_bias_slice = lax.dynamic_slice_in_dim(attn_bias,
(lax.axis_index(axis_name) - idx) q_block_idx = lax.axis_index(axis_name)
k_block_idx = (lax.axis_index(axis_name) - idx) q_chunk_idx_start = q_block_idx * (block_size // query_chunk_size)
k_chunk_idx_start = k_block_idx * (block_size // key_chunk_size)
dq, dk, dv = _blockwise_attention_bwd(q, k, v, g, (dq, dk, dv, output, denominator, max_score),
q_chunk_idx_start, k_chunk_idx_start, bias=attn_bias_slice, **blockwise_kwargs)
k, v, dk, dv = map(lambda x: lax.ppermute(x, axis_name, perm=[(i,
(i + 1) return (dq, dk, dv, k, v), None
dq, dk, dv = dq.astype(q.dtype), dk.astype(k.dtype), dv.astype(v.dtype)
return dq, dk, dv, None
[…]
def ring_attention(q, k, v, attn_bias, axis_name, float32_logits, blockwise_kwargs):
y, _ = _ring_attention_fwd(q, k, v, attn_bias, axis_name, float32_logits, blockwise_kwargs)
return y
[…]
Appendix B Experiment Details
B.1 Evaluation of context length
In the experimental results presented in Section 5.1, we used fully sharded tensor parallelism (FSDP) to partition the model across GPUs or TPU devices. Our evaluation focused on determining the maximum achievable sequence length in commonly used FSDP training scenarios. For TPUs, we utilized its default training configuration, which involved performing matmul operations in bfloat16 format with weight accumulation in float32. On the other hand, for GPUs, we adopted the default setup, where all operations were performed in float32.
B.2 Evaluation of MFU
In the evaluation presented in Section 5.2. The batch size in tokens is 2 million per batch on GPU and 4 million per batch on TPU. The training was conducted using FSDP Facebook [2023] with Jax SPMD. For gradient checkpointing [Chen et al., 2016], we used nothing_saveable as checkpointing policies for attention and feedforward network (FFN). For more details, please refer to Jax documentation.
B.3 Evaluation on line retrieval
In the evaluation presented in Section 5.4, we finetuned the LLaMA-13B model [Touvron et al., 2023], limiting context length to 512K tokens due to constraints on our cloud compute budget, the training was conducted on 32x A100 80GB Cloud GPUs. We use user-shared conversations gathered from ShareGPT.com with its public APIs for finetuning, following methodologies as outlined in prior works [Chiang et al., 2023, Geng et al., 2023]. ShareGPT is a website where users can share their ChatGPT conversations. To ensure data quality, we convert the HTML back to markdown and filter out some inappropriate or low-quality samples, which results in 125K conversations after data cleaning.
Appendix C Inference requirement
We provide the minimal sequence length required to overlap communication with computation during training in Table 2. Ring Attention enables effortless training of context size that scales linearly with the number of devices. While we focus on introducing training as it is more memory demanding than autoregressive inference where the number of query token is one, Ring Attention is applicable to inference too. For example, serving a LLaMa 7B on 32x TPUv5e, the conventional approach is to distribute the model along the attention heads dimension, with each device computing one attention head. Assuming a batch size of 1, this can serve up to a 256K context length due to key-value cache activation size. Ring Attention can allow 32 times larger context by circulating the key-value cache between a ring of devices. To overlap the communication with computation, it needs d2/F >= 2*d2/B, where B/F >=2. With a bandwidth of 186 GB/s and flops of 196 TFLOPs, and assuming an unreasonably high MFU of 40% for this large context, then B/F = 2.4, meaning that Ring Attention allows 32 times larger context for inference without adding overheads.
Appendix D Training FLOPs Scaling of Context Size
Given that our proposed approach unlocks the possibility of training with a context size exceeding 100 million tokens and allows for linear scaling of the context size based on the number of devices, it is essential to understand how the training FLOPs per dataset scale with the context size. While a larger context size results in a higher number of FLOPs, the increased ratio does not scale quadratically because the number of tokens remains fixed. We present these results in Figure 5, which showcases various model sizes and context lengths, representing different computational budgets. The figure shows the ratio of FLOPs for larger context lengths compared to the same model with a shorter 4K context size.
We calculated the per sequence FLOPs using $(24bsh^{2}+4bs^{2}h)n$ where $h$ is model hidden dimension, $b$ is batch size, $s$ is total sequence length, and $n$ is number of layers. The per dataset FLOPs ratio is then given by $((24bs_{2}h^{2}+4b{s_{2}}^{2}h)/(24bs_{1}h^{2}+4b{s_{1}}^{2}h))/(s_{2}/s_{1})=(6h+s_{2})/(6h+s_{1})$, where $s_{2}$ and $s_{1}$ are new and old context lengths. Model sizes and their hidden dimensions are as follows: LLaMA-7B (4096), LLaMA-13B (5140), LLaMA-33B (7168), LLaMA-65B (8192), GPT3-175B (12288), and 1TB (36864). These model configurations are from LLaMA [Touvron et al., 2023] and GPT-3 [Brown et al., 2020] papers, except the 1TB model size and dimension were defined by us.